Part A What is the change in energy ?E of the hydrogen atom as the electron...

Part A

What is the change in energy ?E of the hydrogen atom as the electron makes the transition from the n=3 energy level to the n=1 energy level? Express your answer numerically in electron volts.

Part B

When an atom makes a transition from a higher energy level to a lower one, a photon is released. What is the wavelength of the photon that is emitted from the atom during the transition from n=3 to n=1? Express your answer numerically in angstroms.

Part C

When the hydrogen atom makes the transition from the n=2 to the n=1 energy level, it emits a photon. This photon can be absorbed by to a singly ionized helium atom (He+) in its second energy level (ninitial=2), causing it to move to a higher energy level. To what final energy level number nfinal can the photon promote the helium ion?

Note that He+ has two protons in its nucleus and a single electron.

Express your answer as an integer.
Answer #1

A) The energy of a photon is related to the wavelength by, hc に10.26×10-7 A hc E=-=-12.1 eV = (1.097×10 m/s)(F-32 λ = 1030 A